Cody was selling drawings for $5 each. For every 8 drawings he sold, he had to use $4 in art supplies to make them. How much profit would he make if he sold 8 drawings?
Answer by ikleyn(52781)

.

His profit is  8*5 - 4 = 40 - 4 = 36 dollars.    ANSWER

Solved, answered and explained.